# Only Send What You Need: Learning to Communicate Efficiently in Federated Multilingual Machine Translation

### v1

@@ Appendix G FL Experiments Without Using Pre-trained Knowledge
| 16.5 | 6.3 | 11.4 | 21.3 | 20.5 | 19.0 | 20.3 | ||
| 16.0 | 6.9 | 11.5 | 20.8 | 21.0 | 19.7 | 20.5 | ||
| MetaSend (ours) | 17.9 | 7.0 | 12.5 | 20.9 | 21.7 | 18.6 | 20.4 | |
| MetaSend (ours) | 17.5 | 7.5 | 12.5 | 22.7 | 20.8 | 20.4 | 21.3 | |
@@
| 14.6 | 6.6 | 10.6 | 16.3 | 14.9 | 15.0 | 15.4 | ||
| 14.9 | 6.8 | 10.9 | 16.6 | 15.4 | 15.3 | 15.8 | ||
| MetaSend (ours) | 14.9 | 6.9 | 10.9 | 18.5 | 16.6 | 16.3 | 17.1 | |
| MetaSend (ours) | 15.1 | 7.3 | 11.2 | 17.7 | 18.2 | 16.0 | 17.3 | |
@@
| 0.625 | 0.647 | 0.636 | 0.700 | 0.669 | 0.718 | 0.696 | ||
| 0.631 | 0.651 | 0.641 | 0.719 | 0.673 | 0.727 | 0.706 | ||
| MetaSend (ours) | 0.634 | 0.662 | 0.648 | 0.722 | 0.679 | 0.728 | 0.710 | |
| MetaSend (ours) | 0.638 | 0.657 | 0.648 | 0.725 | 0.684 | 0.725 | 0.711 | |
@@
| MetaSend (ours) | 0.620 | 0.647 | 0.634 | 0.676 | 0.640 | 0.653 | 0.656 | |
| MetaSend (ours) | 0.619 | 0.648 | 0.634 | 0.676 | 0.638 | 0.659 | 0.658 | |
Appendix H Supplementary Results for Ablation Study
This section provides additional results to Section 5.3. Figure 12 shows the ablation study of our method on the MTNT and UNMT datasets, examining the impact of using different numbers of neurons in our MAML module. On the other hand, Figure 13 presents the ablation study of our method, investigating the effects of using different batches for the MAML module. We observe that these additional results align with the findings discussed in Section 5.3. Specifically, increasing the resources, such as the number of neurons or batches, for our MAML module generally results in improved performance. Additionally, Table 7 provides specific time measurements for meta-evaluation with different numbers of batch samples. Intuitively, it can be observed that using more samples for meta-learning will result in increased time requirements.

### v2

@@ V-C4 The use of pre-trained models
| DPg | 16.5 | 6.3 | 11.4 | 21.3 | 20.5 | 19.0 | 20.3 | |
| DPl | 16.0 | 6.9 | 11.5 | 20.8 | 21.0 | 19.7 | 20.5 | |
| MetaSendg (ours) | 17.9 | 7.0 | 12.5 | 20.9 | 21.7 | 18.6 | 20.4 | |
| MetaSendl (ours) | 17.5 | 7.5 | 12.5 | 22.7 | 20.8 | 20.4 | 21.3 | |
@@
| DPg | 14.6 | 6.6 | 10.6 | 16.3 | 14.9 | 15.0 | 15.4 | |
| DPl | 14.9 | 6.8 | 10.9 | 16.6 | 15.4 | 15.3 | 15.8 | |
| MetaSendg (ours) | 14.9 | 6.9 | 10.9 | 18.5 | 16.6 | 16.3 | 17.1 | |
| MetaSendl (ours) | 15.1 | 7.3 | 11.2 | 17.7 | 18.2 | 16.0 | 17.3 | |
@@
| DPg | 0.625 | 0.647 | 0.636 | 0.700 | 0.669 | 0.718 | 0.696 | |
| DPl | 0.631 | 0.651 | 0.641 | 0.719 | 0.673 | 0.727 | 0.706 | |
| MetaSendg (ours) | 0.634 | 0.662 | 0.648 | 0.722 | 0.679 | 0.728 | 0.710 | |
| MetaSendl (ours) | 0.638 | 0.657 | 0.648 | 0.725 | 0.684 | 0.725 | 0.711 | |
@@
| MetaSendg (ours) | 0.620 | 0.647 | 0.634 | 0.676 | 0.640 | 0.653 | 0.656 | |
| MetaSendl (ours) | 0.619 | 0.648 | 0.634 | 0.676 | 0.638 | 0.659 | 0.658 | |
V-D Varying Data Availability
V-D1 Insufficient data samples
To mirror the limited data scenario of each client in practical FL, we performed experiments by randomly sampling a small portion of data from MTNT and UNMT datasets as training data. Specifically, we randomly selected 20% of samples from each dataset, resulting in 1k and 10k training samples in each language direction for MTNT and UNMT datasets, respectively, while keeping the validation and test sets at the same size. All other hyperparameters, such as batch size for the NMT engine or MAML optimization, neurons in the MAML module, and learning rate, remain the same as in Section V. Tables IX and X present the translation quality of each method when trained with limited data. Even in scenarios with limited resources, it is evident that our MetaSend methods consistently outperform other baselines and, in some cases, achieve comparable or even slightly better performance compared to PMFL.
| Training | Method | MTNT | UNMT | |||||
| En Fr | En Ja | Avg | En Fr | Ar Es | Ru Zh | Avg | ||
| IID FL | PMFL | 27.3 | 11.5 | 19.4 | 33.6 | 33.2 | 34.4 | 33.7 |
| Non-IID FL | PMFL | 16.6 | 8.8 | 12.7 | 19.4 | 19.6 | 17.7 | 18.9 |
| IID FL | RandSend | 26.1 | 10.4 | 18.3 | 31.5 | 31.3 | 32.1 | 31.6 |
| DPg | 27.3 | 10.9 | 19.1 | 32.6 | 32.7 | 32.3 | 32.5 | |
| DPl | 26.8 | 11.0 | 18.9 | 33.0 | 32.9 | 32.8 | 32.9 | |
| MetaSendg (ours) | 28.0 | 11.8 | 19.9 | 34.3 | 33.0 | 34.1 | 33.8 | |
| MetaSendl (ours) | 27.3 | 12.0 | 19.7 | 34.2 | 33.2 | 34.1 | 33.8 | |
| Non-IID FL | RandSend | 14.9 | 7.4 | 11.1 | 18.2 | 17.4 | 16.4 | 17.3 |
| DPg | 15.3 | 7.9 | 11.6 | 18.4 | 19.5 | 16.1 | 18.0 | |
| DPl | 15.0 | 7.3 | 11.2 | 18.9 | 19.0 | 16.4 | 18.1 | |
| MetaSendg (ours) | 15.8 | 7.9 | 11.9 | 19.5 | 19.5 | 17.5 | 18.8 | |
| MetaSendl (ours) | 16.0 | 7.8 | 11.9 | 19.4 | 19.7 | 17.7 | 18.9 | |
| Training | Method | MTNT | UNMT | |||||
| En Fr | En Ja | Avg | En Fr | Ar Es | Ru Zh | Avg | ||
| IID FL | PMFL | 0.687 | 0.683 | 0.685 | 0.750 | 0.736 | 0.745 | 0.744 |
| Non-IID FL | PMFL | 0.640 | 0.661 | 0.651 | 0.677 | 0.664 | 0.671 | 0.671 |
| IID FL | RandSend | 0.666 | 0.655 | 0.661 | 0.733 | 0.719 | 0.727 | 0.726 |
| DPg | 0.701 | 0.678 | 0.690 | 0.746 | 0.727 | 0.739 | 0.737 | |
| DPl | 0.691 | 0.680 | 0.686 | 0.741 | 0.729 | 0.733 | 0.734 | |
| MetaSendg (ours) | 0.711 | 0.680 | 0.696 | 0.752 | 0.736 | 0.743 | 0.744 | |
| MetaSendl (ours) | 0.687 | 0.684 | 0.685 | 0.750 | 0.736 | 0.747 | 0.743 | |
| DPg | 0.633 | 0.654 | 0.644 | 0.673 | 0.649 | 0.659 | 0.660 | |
| DPl | 0.631 | 0.659 | 0.645 | 0.680 | 0.653 | 0.678 | 0.670 | |
| MetaSendg (ours) | 0.640 | 0.670 | 0.655 | 0.659 | 0.670 | 0.711 | 0.680 | |
| MetaSendl (ours) | 0.638 | 0.670 | 0.654 | 0.685 | 0.688 | 0.685 | 0.686 | |
V-D2 Low-resource language scenarios
In addition to scenarios with sparse data samples for every client, we also consider the case of data imbalance, where only one of the clients is designated as low-resource. Specifically, we randomly select 20% of the samples from the En Ja language pair in the MTNT dataset and 20% from the Ru Zh language pair in the UNMT dataset, while maintaining the full training samples for other clients and keeping the validation and test sets unchanged. This selection creates a scenario where a specific language pair (En Ja and Ru Zh) becomes low-resource. Table XI presents the results under this setting. It is evident that the NMT model achieves limited performance for the low-resource client, and the overall performance is also negatively impacted due to this client’s constraints, as compared to the full results for PMFL shown in Table I. However, in comparison to the baseline methods in Table XI, our adaptive thresholding approach delivers the best performance. This underscores the effectiveness of our adaptive thresholding approach in mitigating the challenges posed by low-resource clients, resulting in better overall performance.
| Method | MTNT | UNMT | |||||
| En Fr | En Ja | Avg | En Fr | Ar Es | Ru Zh | Avg | |
| PMFL | 23.1 | 15.6 | 19.4 | 27.9 | 28.4 | 22.9 | 26.4 |
| RandSend | 14.9 | 9.9 | 12.4 | 21.9 | 20.6 | 15.7 | 19.4 |
| DPg | 17.3 | 11.6 | 14.5 | 25.4 | 23.9 | 17.3 | 22.2 |
| DPl | 18.1 | 12.5 | 15.3 | 25.8 | 24.2 | 17.8 | 22.6 |
| MetaSendg (ours) | 19.7 | 13.9 | 16.8 | 25.8 | 23.9 | 17.5 | 22.4 |
| MetaSendl (ours) | 20.1 | 13.6 | 16.9 | 26.1 | 24.7 | 18.0 | 22.9 |
V-E Compatibility with other Pruning Methods
We have shown MetaSend’s ability to learn adaptive thresholds for identifying relevant parameters for communication across various scenarios. To further explore its potential, we examined its integration with general model pruning methods [50, 51, 52]. Given the substantial number of parameters and complex model blocks in language models, directly implementing the additional computations proposed in [50, 51] could significantly increase client computational overhead. Instead, we adapt their concepts by introducing an auxiliary multi-layer perceptron (MLP) layer at the output of the language model to identify redundant parameters, providing further refinement to the parameters pruned by MetaSend. Similar to the importance concept introduced in [51], the MLP takes the magnitudes of the remaining tensors as input and determines whether they should be further pruned. Acting as a second filter, this MLP predicts which parameters are redundant and can potentially be removed to enhance communication efficiency.
Table XII presents the results of integrating our method with the MLP-based pruning approach on the MTNT and UNMT datasets. With this second filtering step, the MLP-based pruning method further reduces parameter transmissions, although the impact on translation performance varies. For MetaSendg, it has no noticeable impact on translation quality, whereas for MetaSendl, it achieves greater parameter savings at the cost of a slight reduction in translation performance. Overall, these results demonstrate the compatibility of our method with other pruning techniques to further optimize communication efficiency, while also highlighting that MetaSend itself effectively identifies and prioritizes the most relevant parameters.
| Training | Method | MTNT | UNMT | ||
| Translation Performance | Tensor Saving | Translation Performance | Tensor Saving | ||
| IID FL | MetaSendg | 24.4 | 61.5% | 41.0 | 59.2% |
| MetaSendg + MLP | 24.3 | +5.4% | 41.1 | +7.9% | |
| MetaSendl | 25.0 | 41.9% | 41.1 | 44.5% | |
| MetaSendl + MLP | 24.7 | +10.0% | 40.8 | +9.7% | |
| Non-IID FL | MetaSendg | 16.8 | 62.3% | 30.3 | 58.3% |
| MetaSendg + MLP | 16.8 | +6.9% | 30.5 | +8.2% | |
| MetaSendl | 16.9 | 42.7% | 31.3 | 43.9% | |
| MetaSendl + MLP | 16.4 | +9.2% | 30.9 | +11.3% | |
VI Conclusion
We conducted the first analysis on the efficiency of federated multilingual NMT. To address the practical challenges that arise in this setup, we proposed MetaSend, which selects tensors for transmission that are most critical to the NMT. Our observation shows that designing a dynamic sending threshold to control the transmitted parameters in each client is crucial during FL communication. By adaptively learning the sending threshold in each FL round based on meta-learning, we saw that our designed MetaSend not only improves communication efficiency, but also effectively captures the NMT threshold for sending. Extensive experiments on two datasets showed that MetaSend outperforms existing baselines regarding machine translation quality and significantly reduces communication costs in FL, confirming its advantage in practical federated NMT settings.